OpenStudy (anonymous):

Which is the best strategy to solve this problem? Mary-Ann is making gift baskets for her friends. She has 32 candles and 64 lotion bottles. She splits the candles and lotion bottles equally among 8 baskets. How many items were in each basket?

OpenStudy (anonymous):

A. Draw a diagram. Draw 8 circles to represent baskets. Think 32 ÷ 8 = 4, so draw 4 candles in each circle. Think 64 ÷ 8 = 8, so draw 8 dots in each circle. Count up the objects in a circle to find the number of items in each basket.

OpenStudy (anonymous):

B. Guess and test. Guess that the number of items in each basket is 9. Multiply 8 × 9. Ask yourself if this is the same as the sum of 32 + 64? If not, revise your guess, and test your answer again.

OpenStudy (anonymous):

C. Write a number sentence. (32 + 64) × 8 =

OpenStudy (alexishope47):

I would guess A.

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

Will you be able to help me?